OpenStudy (anonymous):

A propeller revolves 468 times in 52 s. The number of revolutions varies directly with the amount of time the propeller is revolving. How long will it take the propeller to revolve 756 times? s

OpenStudy (anonymous):

FAN MEDAL

OpenStudy (danjs):

\[\frac{ 468 rev }{ 52 s }\]

OpenStudy (anonymous):

I THINK 9

OpenStudy (anonymous):

\[756\div9=\]

OpenStudy (danjs):

\[\frac{ 468 rev }{ 52s } = \frac{ 756 rev }{ x \sec }\]

OpenStudy (anonymous):

hey sis:)

OpenStudy (danjs):

x = 84 seconds

OpenStudy (anonymous):

The answer is 84. I got that a different way though.